LUCIFER Recalibration after Installation on the Telescope

Executive Summary Block
DAYTIME Tests . . . . . . . . . . . . .
  • LUCI1 re-installation on telescope. Done
  • General function tests Done
  • Internal Focus determination: Done
  • MOS tests Done
  • General reproducibility tests Done
  • Spectral Alignment Check: Done
  • FITS header check Done
  • Field centering and Pupil alignment Done
  • Initial Grating Tilt (Lambda-V) Tests Done
  • Initial lambda-to-voltage for zjSpec Done
  • FlexComp - Calibrate Done
  • FlexComp - Verify Done
  • LUCI Aux cryostat test-fit Fit OK, transfer problem
NIGHTTIME Tests . . . . . . . . . . .
  • Twilight Flats Data taken in J,H,K,BrG,H2,FeII,Pbeta
  • Orientation on sky Done
  • Image scale, all cameras: Done for N3.75
  • Find Center of Rotation Done
  • Calibration of AGw transform Done
  • Field Aberration Measurement
  • Verification of AGw transform Done
  • PM: Photometric ZPs Partial(J,K only)
  • PM: Field illumination H data taken in poor conditions
  • PM: Specphot tests
  • Ghosts, scattered light Done?
  • Run a complete science observation Done
  • LUCI-to-AGw Focus Measurement Accepted as is
  • Collimation model Accepted as is
  • Pointing model Update if possible
  • Run a science observation Done

To get Luci1 science ready for the science-block starting March 12 2012

Aim: To have all data required been taken AND analyzed to give partners a as best as possible starter for the obs.

Based on Commissioning report Lucifer1v1.1, April 2010, in particular section 7.2 "nighttime tests", p. 41 - 79.

Duration and pre-requisites are estimated from report.

Separated into day-time and night-time activities

  • WS: Walter Seifert
  • JH: Jochen Heidt
  • PB: Peter Buschkamp
  • DT: Dave Thompson
  • LBTO: Unspecified LBTO personnel


On-telescope day-time (closed dome) calibrations, daily schedule:

Monday 27 Feb 2012

LUCI1 re-installation on telescope. Done 28.Feb.2012 Duration 1.0d
This is NOT the official LUCI transfer procedure !!!

Responsible: LBTO personnel
Comments: Should take a half day but a full day is scheduled.

  • Stop LUCI SW
  • Shut down Sun_V880
  • Move LUCI to telescope
  • Move Sun_V880 to Computer Room B
  • Re-cable fibers for telescope installation
  • Start Sun_V880
  • Start LUCI SW
  • Check communication with instrument
  • Wait for LUCI to return to thermal equilibrium - Detector stable @77K
  • Switch on ROE, wait 2hr

Tuesday 28 Feb 2012

General function tests Done 29.Feb.2012 Duration: 4h
Responsible: WS
Telescope: @Zenith, Rotator enabled, Lights can be ON
Comments: Use ONLY engineering panels, Main Instrument GUI only AFTER Peter has checked the MOS; start with rotator angle 341deg, repeat for 2-3 other rotator angles

> Basic instrument function tests (from WS email Aug 22, 2011):
  • Stop SW / Re-start SW (optional, restarted in move to telescope)
  • Repeat at ROT ~ 341, 221, 101 degrees:
    • Camera Wheel: Init; move several times between N1.80 & N3.75 camera
    • FW 1 & 2: init, move both wheels front/backwards movements
    • Grating Unit: init, select all gratings/mirror
    • Grating Tilt: set voltages 0V, +-1V
    • Detector focus: move steps to reach neglim & poslim, note # of steps
    • PV: init and move in/out
  • Check that all calib lamps work?

Notes from WS on functional checkout:
------------------------------------------
28.02.

ROT=341
o CameraWheel
   +-737 steps: ok, but at the beginning it needed sometimes the repeat steps
       now no more after being close to thermal balance
o GratUnit
   Selection: ok
   GratTilt
   G210 +0.07/+4.72/-4.82 ok  0.00/+4.45/-4.45 ok
   G200 -0.55/+4.39/-4.53 ok  0.00/+4.00/-4.10 ok
   G150 -1.21/+2.61/-4.65 ok  0.00/+2.20/-4.30 ok
o Focus
   +-630steps, ok & stable, both limit switches fine
o PV
   ok
o FW1 & FW2
   both ok, very rarely needed one repeat step, but no error at all

-> Direct Image with N3.75 & N1.80 looks strange: AGW not in park & M3 reflecting dome lights...
       luci1_20120228_0001+2.fits

o OpticSieve in place
   N3.75 #0003, 0005, 0007     all positions within <1pix
   N1.80 #0004, 0006, 0008     all positions within <1pix

ROT=251
o Image
   N1.80 #0009    flex<2pix
   N3.75 #0010    flex<2pix
o GratUnit
   Selection: ok
   Tilt
   G210 +-3.8V closed loop
   G200 +-3.8V closed loop
   G150 +1.6/-3.8 closed loop
o FW1 & FW2
   1 ok
   2: issue when moving from pos2 (blind) to 1(clear): needs re-init
oCamWheel
   at first ok, but then issue with the steps, needs several times re-init
   steps changed to +-740 & +-15
   but N1.80 position is 'moving' -> changed to 738steps, but no more time left today
   ok
o PV
   ok
o Focus
   +-630 steps, ok

------------------------------------------
29.02.

still ROT=251
o FW2
   now ok w/o any change ?
o CamWheel
   steps changed to +744 7 -740 steps
   stable
   N3.75 position identical to before
   N1.80 position stable, but changed by about 30pix in y

o Obs @ LBTO, no change of ROT possible

o PupilAlignment
   Field with N3.75 camera has been aligned before
   pupil before alignment: luci1_20120229_0001.fits
   Solution of linear system:
   center of pupil stop: 915,1151; center of M2 image: 881,1198
   dx=+35pix, dy=-50pix
   Mirror    M1(y)    M2(x)
   FM1        -7000    -4000 #0002
   FM4        -2000    -3000 #0003
   iteration needed as center of M2 image: 931,1150
   dx=-15pix, dy=0pix
   FM1 M2 +2000
   FM4 M2 -1000        #0004
   -> ok, dx=dy=0 +- few pix

   Now check image
   N3.75 with optic sieve #0005
   N1.80 w optsieve #0006

ROT=161
o CamWheel
   ok
   +744 -741 steps
   all within +-1pix
o Focus:
   ok, +-630steps
o GratUnit
   ok
o PV:
   ok
o FW1 & 2
   ok, both
o Mask exchange
   ok (#1 & #2)

ROT=71deg
o CamWheel
   ok
   +744 -740 steps
   all within +-2pix
o Focus:
   ok, +-630steps
o GratUnit
   ok
o PV:
   ok
o FW1 & 2
   ok, both
o Mask exchange
   ok (#2 & #1 & #2)

Internal Focus determination: Done 29.Feb.2012 Duration: 0h
Responsible: WS
Telescope: @Zenith, Rotator @ 341, Lights can be ON
LUCI: Optic Sieve mask
Done as part of the general function tests, listed here to record results

> Detector stage focus calibration:
  • See ~engineer/TemplateScripts/focusN180.txt and focusN375.txt
  • adjust step range if needed
  • Insert best focus values for both cameras into detector unit GUI config
  • Check tilt of detector w.r.t. focal plane in these data?

Best Focus values:
  • N1.80
  • N3.75

MOS tests Done 29.Feb.2012 Duration: 4h
Responsible: PB
Telescope: @Zenith, Rotator enabled, Lights can be ON

> Basic MOS checkout:
Tests not detailed. Peter's steps to complete the MOS checkout should be documented for future reference.

 #############################################################################
 # Once the MOS tests have been completed (and passed!) and both Walter and  #
 # Peter have given their OKs to continue, you can use the main instrument   #
 # GUI for all remaining tests.  Be sure to start the telescope service in   #
 # the LUCI SW control GUI manually.                                         #
 #############################################################################

General reproducibility tests Done 29.Feb.2012 Duration: ?h
Responsible: JH, DT
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, Spectro Sieve mask.
Comments:
  • move ~5x between the corresponding units
  • Repeat at 3-4 rotator angles (90deg separation)
  • Take image after each move
  • Measure reproducibility

Reproducibility tests:
  • Grating Unit
    • Use arclamp (or arclamp + flatlamp to see continuum?)
    • only moves grating unit between measurements
    • All gratings + N3.75 combos, see table below
    • Closed tilt loop on grating (request nominal CWL)
    • Move between mirror <-> grating 5x
    • take image at each position
  • Cameras
    • Use flatfield lamp or thermal background
    • Only moves camera wheel between measurements
    • Grating Unit fixed on mirror position
    • Move between N1.80 <-> N3.75 5x
    • take image at each position
  • Calculate rms positioning for each set of data

RESULTS
Camera wheel reproducibility test results
   RMS in pixels at ROT=71, 161, 251, 341; Nsamp=5 for all configurations

   CamPos    X___071___Y  X___161___Y  X___251___Y  X___341___Y   Datafile
   N3.75     0.077 0.590  0.128 0.497  0.123 0.525  0.056 0.432   N375_cam.dat
   N1.80     5.367 15.71  7.167 15.63  5.424 32.02  4.425 25.57   N180_cam.dat

   Camera test repeated at ROT=341 four times, so Nsamp=20
   CamPos    X___341___Y   Datafile
   N3.75     0.050 0.315   N375_at341_cam.dat
   N1.80     3.676 21.15   N180_at341_cam.dat

Grating wheel reproducibility test results
   RMS in pixels at ROT=71, 161, 251, 341; Nsamp=5 for all configurations

   GratPos   X___071___Y  X___161___Y  X___251___Y  X___341___Y   Datafile
   Mirror    0.176 0.020  0.395 0.041  0.387 0.019  0.279 0.055   Mirror_grat.dat
   210_zJHK  0.213 0.009  0.550 0.038  0.561 0.041  0.717 0.025   210_grat.dat
   200_H+K   0.436 0.089  0.741 0.033  0.514 0.032  0.767 0.038   200_grat.dat
   150_Ks    0.413 0.018  0.574 0.029  0.459 0.038  0.511 0.027   150_grat.dat

Datafiles attached to this page as Repro_dat.tar.gz

Note from WS: Incorrect motor step value in configuration.  Fixed, and now N1.8 repositions well below 1 pixel rms. 

Notes from the data collection:
Data taken for these tests @ ROT=71
luci_20120229_CamRep_0001-0010 (even=N3.75, odd=N1.80)

All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
luci_20120229_GratRep_* (pairs are "off" then "on" for the Ar lamp)
Mirror 0001 0005 0009 0013,14 19,20
210_zJHK 0002 0006 0010 0015,16 21,22
200_H+K 0003 0007 0011 0017,18 23,25
150_Ks 0004 0008 0012 __0001,2 24(on)

Data taken for these tests @ ROT=161
luci_20120229_CamRep_0011-0025

Got error message while executing the script "error setting camera position" when moving from N3.75 to N1.80 camera.

Initialized camera-wheel => OK

Use only CamRep_0016-0025 for analysis (even=N3.75, odd=N1.80)

All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
luci_20120229_GratRep_*
Mirror 0026 0030 0034 0038 0042
210_zJHK 0027 0031 0035 0039 0043
200_H+K 0028 00032 0036 0040 0044
150_Ks 0029 0033 0037 0041 0045

Data taken for these tests @ ROT=251
luci_20120229_CamRep_0026-0035 (even=N3.75, odd=N1.80)

All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
luci_20120229_GratRep_*
Mirror 0046 0050 0054 0058 0062
210_zJHK 0047 0051 0055 0059 0063
200_H+K 0048 00052 0056 0060 0064
150_Ks 0049 0053 0057 0061 0065

Data taken for these tests @ ROT=341
luci_20120229_CamRep_0036-0049

Got error message while executing the script "error setting camera position" when moving from N1.80 to N3.75 camera.

Initialized camera-wheel => OK

Use only CamRep_0040-0049 for analysis (even=N3.75, odd=N1.80)

As a check for camera-failures, the camera-repro scripts were repeated 3 times.

=> no further error messages. Images CamRep_0050-0079

All @ clear+K, 2.2um CWL, mirror=4s, gratings=30s, Ar lamp on
luci_20120229_GratRep_*
Mirror 0066 0070 0074 0078 0082
210_zJHK 0067 0071 0075 0079 0083
200_H+K 0068 00072 0076 0080 0084
150_Ks 0069 0073 0077 0081 0085

The images with the N1.80 camera show strong defocus/coma in the upper/lower part of the image. Images have been taken with the OptikSieve, J/OH_1190, mirror/halo1 (on/off for badpixel correction).

N1.80: luci1_20120229_CamRep_Test_0005-0006

N3.75: luci1_20120229_CamRep_Test_0007-0008

Images Test_0001-4 are not useful (saturated)

Notes from WS post-testing check of N1.8 camera:
o As Cam-Repro gave unacceptable result last night, re-done
   mistake in steps towards N1.80 camera, value in data base
   was -742steps, but should be -740!!! (see above)
   -> corrected and run for 4 rotator angles, 5 positions
       for each angle; no images saved, measured directly on
       GEIRS display

   Result: (given are max. +- deviations [pix])
   ROT        N3.75        N1.80
   deg        x    y        x    y
   --------------------------
   341        0    0        0    1
   251        0   0        1    1
   161        0    1        0    0
    71        1    0        0    1
   341        0    0        0    0

   This result is fully acceptable. Repeat of the cam repro scripts seesm not necessary


o As N1.80 camera has been always in beam when chaning rot. angle,
   a first guess of the flexure can be made. Given are deviations
   between the angles in pix.

   ROT        N3.75        N1.80
   deg        dx    dy        dx    dy
   --------------------------
   341        0    0        0    0
   251        +8  +8        +3    +4
   161        -7    -8        -4    -2
    71        -6    -5        -2    -3
   341        +6    +5        +3    +3

   This looks not bad ;-)

o TempSensor Camera (#5) shows sometimes different temperature and gives warning.
   As this is not a real change, the warning has been switched of in the supervisor config.

Spectral Alignment Check: Done Duration: 0h
Responsible: DT, JH
Note: uses data from Grating Unit reproducibility test above?

### Would this be better done with a flat lamp instead of an arc lamp?

  • For each grating+camera combo in table below:
    • Make sure grating tilt stabilization is ON
    • Set to grating to nominal CWL, take dispersed image
    • Set to mirror, take image
    • Only 1 pair of exposures needed for each combination
  • Measure spectrum tilt along the rows
  • Measure vertical (y) displacement between image and spectrum

Define:
  • Ytilt = Ymax - Ymin along dispersed spectrum near Y~1024
  • Yoffset = Yimg - Yspec for a centered pinhole @ X~1024

Data: luci.20120229.*
Spectral Alignment Ytilt Yoffset Img#
N1.80 + 210_zJHK 2.40 4.07 0047
N1.80 + 200_H+K 0.57 -9.65 0048
N1.80 + 150_Ks 2.15 3.02 0049

Data extracted on the visible continuum, covering half or more of the width of the detector. Data for the N3.75 have the same angular tilt but the Yoffset will be larger by the ratio of the pixel scales (~2.1).

FITS header check Done Duration: 0h (offline)
Responsible: JH, DT
  • Check the FITS headers for any 'strangeness' and note any other issues

Need to update (for N3.75 camera):
  • PIXSCALE = 0.11782
  • CRPIX1 = 1010.58
  • CRPIX1 = 1047.14
  • SECPIX1 = -0.11782
  • SECPIX2 = 0.11782
  • CDELT1 = 3.272778e-05
  • CDELT2 = 3.272778e-05
  • CD matrix as appropriate
Note apparently no status variable for flexure compensation (ON/OFF)!
  • M4M*POS not an indicator (FC ON but these numbers are not changing)
  • Mirror 4 (M4) not the unit that does flexure compensation?

Wednesday 29 Feb 2012

Field centering and Pupil alignment Done 29.Feb.2012 Duration: 5h
Responsible: WS
Telescope: @Zenith, Rotator enabled
LUCI: Imaging, N3.75, K band

> Field centering procedure using FM4
  • Take an image and save it
  • Measure gaps at edges of image
  • FM4: move the mirror in steps to center, notes:
    • Enlarge CompensationMirrorGUI to full screen, use right side of it.
    • M1 -> y, M2-> x; + x steps -> moves image to the right, +y to the top
    • 100 steps corresponds to approximately to 1pix
  • Take images with both cameras, verify centering

> Pupil alignment procedure:
  • Take image to show present alignment of the pupil
  • Adjust FM1 + FM4 (WS: document procedure for future reference)?
  • Repeat until alignment is acceptable

Initial Grating Tilt (Lambda-V) Tests Done Duration: 2d
Responsible: JH, DT
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, LS0.75_450um slit, grating tilt loop closed.

Spectra: grat*
Throughslit-images: ls*
Done for N1.80-camera only

Test table:
Camera Filter Grating Mask Lamp Voltages to test
N1.80 HKspec 200_H+K LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear z 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear J 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear H 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear K 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear Ks 150_Ks LS0.75_450um Ar/Xe/Ne 3 different

First results (measured position of line at the detector in x, input was lambda of line):
If all is OK, lines should always be at the same position within a few pixel

Grat mue x-pos mue x-pos mue x-pos
150Ks 2.0323 1022 2.1714 984 2.3572 928
200H+K 1.5422 955 2.0622 940 2.3572 922
210z 0.8822 1227 0.9516 1088    
210J 1.1691 1138 1.2443 1023 1.2627 996
210H 1.5422 1139 1.6441 1023 1.6733 994
210K 2.1339 1301 2.2046 1231 2.3572 1107

=> The lines shift. Shift is roughly a function of resolution
=> For increasing lambda lines shift towards shorter x-pos ==> "Slope" of tilt (lambda-to-Voltage) is less steep as it was before,
i.e. for a given lambda-range less Voltage coverage is required.

*Not clear where this comes from. Best solution is a complete re-calibration of all grating/order combinations. Tbd as part of the on-sky calibration (e.g. in the afternoon before observation start and/or during periods of poor weather.) *
NO showstopper for on-sky calib.

* Recalibration was done for all gratings and settings during periods of bad weather/technical problems with the AGw3 on March 6 and 7. The wavelength selected at a given observation will correspond to the x-center of the detector (x = 1024). The deviation from x = 1024 across the entire wavelength offered by a grating/setting is always 5pixel at most and less than 2 pixel close to the nominal central wavelength.

Initial lambda-to-voltage measurements for zjSpec Done Duration: ?h
Responsible: JH, DT
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, LS0.75_450um slit.
Note: Approximate center wavelength 1.11um, the midpoint of 0.9um from the dichroic and 1.32um for the red end of J. Initial voltage should then be approximately the same as for HKspec to center 2.22um on the detector.

> Listed separately as this is lower priority than other modes.

Camera Filter Grating Mask Lamp Voltages to test
N1.80 zJspec 200_H+K LS0.75_450um Ar  
N3.75 zJspec 200_H+K LS0.75_450um Ar  

Thursday 01 Mar 2012

Flexure Compensation - calibration data Done 02.Mar.2012 Duration: 8h
Responsible: DT, JH
Telescope: EL and Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam
Scripts located in ~observer/LBTO/FlexComp_Calib/

> Flexure compensation calibration:
  • Exposure times need to be checked (4s for img, 20s for spec)!
  • MANUALLY TURN FLEXURE COMPENSATION OFF to measure!
  • Best sequence:
    • one config
    • all ROT positions (script does this, ROT starts/ends @ -88d)
    • 3 EL's: 90, 55, 20 degrees (order does not matter)
    • next config
  • Use AR lamp for the grating measurements!
  • Process data following ~/LBTO/Flexures/FlexureCheck_Luci1.pdf
  • Select same source in all three datasets for one config!
  • Rename object_0001.tbl like: N180_G210_E90.tbl (see rename.sh)
  • Send renamed tables to Walter

Camera Filter Grating Mask Lamp Script Name Status
N3.75 clear K Mirror Spectro Sieve Thermal FC_N375_mirror.cal Done 01.Mar.2012
N1.80 clear K Mirror Spectro Sieve Thermal FC_N180_mirror.cal Done 01.Mar.2012

Turn on Ar lamp manually!

Camera Filter Grating Mask Lamp Script Name Status
N1.80 HKspec 200_H+K Spectro Sieve Ar FC_N180_200HK.cal Done 01.Mar.2012
N1.80 clear J 210_zJHK Spectro Sieve Ar FC_N180_210zJHK.cal Done 01.Mar.2012
N1.80 clear Ks 150_Ks Spectro Sieve Ar FC_N180_150Ks.cal Done 02.Mar.2012
             
N3.75 clear Ks 150_Ks Spectro Sieve Ar FC_N375_150Ks.cal Done 02.Mar.2012
N3.75 clear J 210_zJHK Spectro Sieve Ar FC_N375_210zJHK.cal Done 02.Mar.2012
N3.75 HKspec 200_H+K Spectro Sieve Ar FC_N375_200HK.cal Done 02.Mar.2012

Data: /data/luci/20120301/luci.20120301.
Scripts: /home/observer/LBTO/FlexComp_Calib
Processing: /home/observer/LBTO/Flexures (see cp.sh and rename.sh)

Flexure Compensation set to OFF when collecting these data!
Script: FC_N180_mirror.cal
EL  Files         midas.tbl            plot.ps 
90  0063 - 0092   N180_MIRROR_E90.tbl  N180_MIRROR_E90.ps
55  0093 - 0122   N180_MIRROR_E55.tbl  N180_MIRROR_E55.ps
20  0123 - 0152   N180_MIRROR_E20.tbl  N180_MIRROR_E20.ps

Script: FC_N375_mirror.cal
EL  Files         midas.tbl            plot.ps 
20  0153 - 0182   N375_MIRROR_E20.tbl  N375_MIRROR_E20.ps
55  0183 - 0212   N375_MIRROR_E55.tbl  N375_MIRROR_E55.ps
90  0213 - 0242   N375_MIRROR_E90.tbl  N375_MIRROR_E90.ps

Script: FC_N180_210zJHK.cal
EL  Files         midas.tbl            plot.ps 
90  0243 - 0272   N180_G210_E90.tbl    N180_G210_E90.ps
55  0273 - 0302   N180_G210_E55.tbl    N180_G210_E55.ps
20  0303 - 0332   N180_G210_E20.tbl    N180_G210_E20.ps

Script: FC_N180_200HK.cal
EL  Files         midas.tbl            plot.ps 
20  0333 - 0362   N180_G200_E20.tbl    N180_G200_E20.ps
55  0363 - 0392   N180_G200_E55.tbl    N180_G200_E55.ps
90  0393 - 0422   N180_G200_E90.tbl    N180_G200_E90.ps

20120302 Continue with calibrations...
Data: /data/luci/20120302/luci.20120302.

Script: FC_N375_150Ks.cal
90  0001 - 0030   N375_G150_E90.tbl    N375_G150_E90.ps    0031=lamp off
55  0032 - 0061   N375_G150_E55.tbl    N375_G150_E55.ps
20  0062 - 0091   N375_G150_E20.tbl    N375_G150_E20.ps

Script: FC_N180_150Ks.cal
20  0101 - 0130   N180_G150_E20.tbl    N180_G150_E20.ps    0094=lamp off
55  0131 - 0160   N180_G150_E55.tbl    N180_G150_E55.ps
90  0161 - 0190   N180_G150_E90.tbl    N180_G150_E90.ps

Script: FC_N375_200HK.cal
90  0201 - 0230   N375_G200_E90.tbl    N375_G200_E90.ps    0192=lamp off
55  0231 - 0260   N375_G200_E55.tbl    N375_G200_E55.ps
20  0261 - 0290   N375_G200_E20.tbl    N375_G200_E20.ps

Script: FC_N375_210zJHK.cal  
20  0291 - 0320   N375_G210_E20.tbl    N375_G210_E20.ps    0191=lamp off
55  0321 - 0350   N375_G210_E55.tbl    N375_G210_E55.ps
90  0351 - 0380   N375_G210_E90.tbl    N375_G210_E90.ps

Note from WS: all the new flex comp files are in place and activated (03.Mar.2012)

Friday 02 Mar 2012

Additional Grating tilt tests?
Additional Flexure compensation calibrations (measurements completed, tabulated above)?
Verification of lambda-to-voltage measurements for zjSpec Duration: ?h

Saturday 03 Mar 2012

Flexure Compensation - verification data Done Duration: 8h

Responsible: OKuhn (both JH and DT en route to LBT).
Telescope: EL and Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam (OK for imaging too, same temp as dome)
Scripts located in /home/observer/LBTO/FlexComp_Verify/
Comment: Verification, only a subset needed
Prereq: Flexure calibrations in place, flexure compensation ON!

> Flexure compensation verification:
  • MANUALLY TURN FLEXURE COMPENSATION ON to verify FC!
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 1 deg/sec
  • Set EL, send LUCI rotator to -69 degrees to start
  • Best sequence:
    • one config
    • all ROT positions (script does this, ROT starts/ends @ -69d)
    • 2 EL's: 38 and 72 degrees (mid-points of measurements)
    • next config
  • Process data following ~/LBTO/Flexures/FlexureCheck_Luci1.pdf
  • Copy results files into ~/LBTO/Flexures/FlexComp_Verify
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 5 deg/sec
  • Send resulting data to Walter

Camera Filter Grating Mask Lamp Script Name
N3.75 clear K Mirror Spectro Sieve Thermal FC_N375_mirror.cal

Turn on Ar lamp manually!

Camera Filter Grating Mask Lamp Script Name
N1.80 HKspec 200_H+K Spectro Sieve Ar FC_N180_200HK.cal
N1.80 clear J 210_zJHK Spectro Sieve Ar FC_N180_210zJHK.cal

Other notes: subtract a dark or a lamp-off before processing. Should see <~2 pix residuals.

Data: /data/luci/20120303/luci.20120303.
Scripts: /home/observer/LBTO/FlexComp_Verify
Processing: /home/observer/LBTO/Flexures (see cp.sh and rename.sh)
Results: /home/observer/LBTO/FlexComp_Verify_Mar2012/

Set Rotator max speed to 1 degree/sec.

Flexure Compensation set to ON when collecting these data!
Script: FC_N375_mirror.cal

EL  Files         midas.tbl            plot.ps 

72  0001-  0011   series of 10 4-sec darks
72  0011 - 0038   N375_MIRROR_E72.tbl  N375_MIRROR_E72.ps, CALunit IN (by accident),  dX~4pxl, dY~5pxl 
38  0039 - 0066   N375_MIRROR_E38.tbl  N375_MIRROR_E38.ps, CALunit OUT, dX~4pxl, dY~5pxl

flush detector, take 20-sec dark with blind/blind luci.20120303.0067.fits
CALunit IN, Ar ON

Next set (0024 - 0051) has FC OFF. I had not touched this button on the IM GUI since starting, however, 
although "Set Flexure Comp" shows ON selected, the box above this read
OFF and was grey, not ON and green like it was for the MIRROR scripts. FC appears to have turned off (in the change of configuration?)
Allowed script to finish, after which I manually selected ON for Set Flexure Comp and now it is ON and green throughout script. 

38  303.0068 - 304.0023 !!!FC OFF!!! N180_G200_E38_fcOFF.tbl  N180_G200_E38_fcOFF.ps, dX>~3pxl, dY~4pxl
UT date changed... luci.20120304.0024 ... 
38  0024 - 0051  N180_G200_E38.tbl N180_G200_E38.ps, dX~1.2pxl, dY~1.5pxl
72  0052 - 0079  N180_G200_E72.tbl N180_G200_E72.ps, dX~2.0pxl, dY~1.5pxl

Immediately ran next script, FC_N180_G210, but found that it turned off FC. Aborted first exposure, cntl-c'd script and then set FC ON. 
Re-ran script and verify FC Status is ON. (Changing configuration does apparently turn off FC and it needs to be turned on again manually or through the script).
72  0080 - 0107  N180_G210_E72.tbl N180_G210_E72.ps, dX~ 3 pxl, dY~1.7 pxl  HUGE JUMP in X (~8 pxls) on image 0092. 
38  0108 - 0135  N180_G210_E38.tbl N180_G210_E38.ps, dX~ 2 pxl, dY~1.1 pxl 

Done with planned set of tests. Summary:

  • N375 and mirror: scatter in X,Y centroids is high, dX~4 pxl and dY~5 pxl.
    • Will repeat at least one elevation.
  • N180 and G200: dX~1.2-2.0 pxl, dY~1.5 pxl
  • N180 and G210: dX~2-3 pxl, dY~1-2 pxl.
    • Jump in X by ~8 pxls during image number 0092.

Repeating N375 mirror since scatter in X, Y centroids was high and there was an error message when turning FC on, although it 
appeared to be ON (green background and status = ON) during the tests.

0136 is a new 4-sec dark
Selected FC ON and verify FC is on after script has been started. No error message this time.

38 0137 - 0164 N375_MIRROR_E38_2.tbl N375_MIRROR_E38_2.ps, dX~4 pxl, dY~ 4pxl.
72 0165 - 0192 N375_MIRROR_E72_2.tbl N375_MIRROR_E72_2.ps, dX~4 pxl, dY~ 9pxl (large scatter but not a single point driving it)

Sunday 04 Mar 2012

(Daytime LBTO schedule currently open)

Monday 05 Mar 2012

LUCI Aux cryostat test-fit

On-telescope night-time on-sky calibrations, daily schedule:

Note:
  • PM = Photometric

Twilight Flats : Done for J, H, K, BrG, H2, FeII, Pbeta Duration: 0h (Twilight)
Evenings/Mornings when photometric:
Pre-requisite: clear skies
Twilight Flatfields

20120304: luci.20120304. O2DCR, 1x4.0s, normal*
  • 0195 - 0204 BrGam ~15K, strong (x2) roll off to left edge!
  • 0205 - 0214 H2 ~16K
  • 0215 - 0224 BrGam ~12K
  • 0225 - 0234 H2 ~12K
  • 0235 - 0244 BrGam ~ 8K
  • 0245 - 0254 H2 ~ 7K
  • 0255 - 0264 FeII ~12K (sunset about here)
  • 0265 - 0274 Pbeta ~ 8K
  • 0275 - 0284 BrGam ~ 1K
  • 0285 - 0294 H2 ~ 1K
  • 0295 - 0304 FeII ~ 1K
  • 0305 - 0314 Pbeta ~ 1K

O2DCR, 1x10.0s, normal
  • 0315 - 0324 J ~ 5K
  • 0325 - 0334 J ~ 4K [Correct ptg/coll. Had 10x10s on dither 0335-0350]
  • 0351 - 0360 J ~ 1K

0361 1x120s MER10 blind+blind
  • Strong persistence visible from pointing correction!

Morning twilight flats in K (random10x, on FS33 field) O2DCR, 1x4.0s, normal
  • 0528-0537 K ~ 2K
  • 0538-0547 K ~ 4...6K
  • 0548-0557 K ~ 6...15K

Evening twilight flats in H (random30x) 1x4.0s 20120305 0036 - 0065 ~16K@start to 3K@end

Flats look OK, no problems noted. Flats taken at different illumination levels can show differences of up to ~5% or so. Recommend keeping the twilight flatfield counts below ~15000 ADU if at all possible. Narrowband flats must be started prior to sunset but can be completed after. Broadband flats need to start immediately after sunset or be taken in the last ~30 minutes before sunrise.

20120305: luci.20120305. O2DCR, 1x4.0s, normal*

*0347-0346, Ks, telescope not tracking (just stationary), about 11-14K

Orientation on sky: Done Duration: 2h
Responsible: DT, JH
Pre-requisite: Orientation done before AGw transform calibration!
Pre-requisite: field with good astrometric coords, good seeing (<1")
Note: Can be done on Persson stds during PM night
Note: Will update further after some MOS acquisitions

> Orientation data
  • Configure for imaging with N3.75, Ks
  • Preset to field (Persson standards, any low-lat field)
  • Take image pair, script: /home/observer/LBTO/SciencePrep/imagepair
  • Repeat for ~5 fields
  • Derive PA offset, update LEFTZEROPOINT
  • Update PCSInstrument.conf and RESTART PCS!
  • Repeat observations for ~5 fields
  • Check that orientation is OK

LEFTZEROPOINT updated to -187.814 (from -187.764). Orientation at PAreq=0.00 is -0.029 ± 0.036 degrees (31 measurements).

Note: systematic residuals in the rotator calibration will show as a consistent rotator offset in MOS mask acquisitions that can be "mined" from the event log after science blocks and the rotator ZPs updated. The systematic dX and dY offsets from MOS alignments should also be a measure of the difference between the pointing origin and the mechanical mask center. This should be true for MODS as well as LUCI.

Image scale, all cameras: Done for N3.75 Duration: 0h
Responsible: DT, JH
Uses orientation or photometry data so 0h needed
N3.75 scale expected to change because of lens swap
Pre-requisite: fields with good astrometric coords, good seeing (<1")
Note: Can use the orientation data (scale is a byproduct)
Note: scale measured by WS from sieve mask data: N3.75=0.118"/pix, N1.80=0.249"/pix

  • N3.75 = 0.11782 ± 0.00005 (31 samples, standard geometric WCS fitting in iraf)
  • N1.80 =

Find Center of Rotation Done Duration: 3h
Responsible: DT, JH
Pre-requisite: Any dense field
Use pointing mask to also measure mechanical center of masks

> Rotator center
  • Configure for imaging with N3.75, Ks
  • Preset to any dense star field at PA=0
  • Allow to collimate
  • Take image
  • Preset to same field at PA=180
  • Allow to collimate
  • Take image
  • Also do PA=90, 270
  • Repeat on at least two different fields

> Mean coords of matched stars gives rotator center > Mean coords of reference circle gives mechanical center

Messier67 field at 08:51:00.0 +11:48:30.0 gs=08:51:02.8 +11:51:25.73 PA=270 Im# PA MASKCEN ROTCEN 394 180 1021.73,1033.46 1009.93,1049.02 395 0 1015.80,1030.61 396 90 1016.08,1026.68 1013.13,1050.12 397 270 1025.83,1039.26 398 dark

  • ROTCEN @ 1010.58 1047.14
  • MASKCEN @ 1019.87 1032.50
  • ROTCEN -> MASKCEN +1.09"x -1.73"y (DETXY coords)

The latter offset will be the typical MOS mask alignment offset needed (this is my prediction).

Additional data was taken on 20120305:
  • 0089 - 0092 (0, 180, 90, 270 degree PAs, respectively)
  • 0-180 data (confirms ROTCEN from previous night):
    • ROTCEN from GEOMAP: 1010.36 1048.04
    • ROTCEN from MATCH: 1010.42 1047.99
  • 90-270 data:
    • ROTCEN from GEOMAP: 1013.31 1049.34
    • ROTCEN from MATCH: 1013.22 1049.49 (3 pixel offset within flexure residuals)

ROTCEN will be different in the N1.80 camera, but it is mostly relevant for N3.75 as that is what is used for mask alignment.

Calibration of AGw transform Done Duration: 6h
Responsible: DLMiller, JStorm (DT, JH to take data?)

See: http://wiki.lbto.org/twiki/bin/view/Commissioning/TransformationData

  • 20120305 transform_collect on StoneE_m90 0093-0109
  • 20120305 transform_collect on Stone_Em90_allgood 0110-0124

Field Aberration Measurement
Responsible: DLMiller (DT, JH to take data?)

See: http://wiki.lbto.org/twiki/bin/view/Commissioning/FieldAberrationMeasurementDetails

Verification of AGw transform Done Duration: 1h
Responsible: ?LBTO (DT, JH to take data)

Once transform is in place, take another set of data to verify. Note: a repeated dice-5 pattern should be well-aligned with rows/columns

20120305: luci.20120305. O2DCR, 5x4.0s, integrated*
Nodding along slit, uses Dave's slitdither.sci script, 20 positions
  • 0289 - 0292 acqusition and centering
  • 0293 - 0312 nodding

Results from centroids of the 10 samples at each position
Xcen(rms) Ycen(rms)
1023.420(0.431) 1518.292(0.277)
1021.860(0.468) 501.285(0.251)

So the dither pattern is rotated slightly on the detector (0.09 degrees, in the same sense as the orientation residual). But the image of the slit is also slightly rotated from perfectly vertical and in the same direction (0.1 deg), such that the dithering along the slit looks extremely good.

Center of slit near the two dither positions (avg of 20 frames): 1023.65 1505.00 1021.65 490.00

PM: Total system throughput and photometric ZPs: Partial (only J, K) Duration: 7h
Responsible: DT, JH
In all BB-filters
Pre-requisite: photometric conditions

### Standard with a range of colors observed at a range of airmass to
### derive ZP@airmass=1 and extinction, perhaps also color terms.

> Photometric zero points
  • Preset to series of photometric standards (AIRM 1.0…2.0)
  • Take imagepair through each BB filter (K, Ks, H, J, z)
  • Minimum ~5 samples per filter spread in airmass, 10 is better

20120304 O2DCR 5x4.0s integrated, Script: /home/LBTO/ScienceOps/dice5_once
ID EL J EL K EL J
FS23 84 0463-0467 70 0488-0492 -- ---
FS20 30 0468-0472 -- --- -- ---
FS21 52 0473-0477 42 0493-0497 -- ---
FS33 61 0478-0482 53 0498-0502 41 0523-0527
FS23 71 0483-0487 62 0503-0507 55 0518-0522
FS27 -- --- 81 0508-0512 82 0513-0517

20120305
ID EL H PA PeakCounts
FS106 45 0069-0073 45 Saturated
FS111 58 0074-0078 00 Saturated
FS116 63 0079-0083 158 Saturated
FS120 71 0084-0088 55 Saturated
FS120 51 0125-0129 180 Saturated
FS121 43 0130-0134 00 Saturated
FS 14 49 0135-0139 100 6K/DIT
FS 15 68 0140-0144 95 25K/DIT
FS125 85 0145-0149 225 Saturated

See attached plot (JKzeropoints.jpg: blue=J, green=H, red=K)
  • J = 25.049 - 0.096(AIRM)
  • K = 24.446 - 0.052(AIRM)

  • JHK Zero Points, March 2012:
    JHKzeropoints.jpg

Results are consistent with a higher QE at shorter wavelengths, but the UM does not state an extinction correction nor whether the tabulated values are for airmass 1 or 0 so it is difficult to compare quantitatively. The seeing was so good on night 2 when the H standards were taken that all but two saturated! The unsaturated sources are consistent with expectations from the J and H data.

PM: Field illumination: H data taken in poor conditions Duration: 3h
Responsible: DT, JH
N3.75 camera due to new field lens
Pre-requisite: photometric conditions
3h should be enough to do a couple of filters
See /home/observer/LBTO/SciencePrep/Illumination for generic scripts

> Field Illumination
  • Configure for imaging in N3.75 camera
  • Select filter, target (any Persson standard, other fields?)
  • Test DIT, configure NDIT to allow collimation
  • Center source on detector
  • Run at least two scripts per filter:
    • Ddice5_100as_illum.scr (gets the extreme corners, +/-100")
    • D25x_illum1.scr (covers field, monitors extinction)
  • Repeat for all filters (eventually) BB and NB.

Ghosts, scattered light: Done? Duration: 2h
Responsible: DT, JH
Not a must but nice add-on.
Pre-requisite: bright star
Bad weather filler

> Ghost check
  • Configure for imaging, J band, N3.75 camera
  • Point to bright star (any in M5 catalog, pref. not K giants)
  • Offset to put the source in the center of the top-left quadrant
  • Take a 7x4.0s exposure
  • Offset DETXY dy=-120" to place in bottom-left quadrant
  • Take a 7x4.0s exposure
  • Difference to check for ghosting, scattered light

Ghosts showed up in some of the photometry data. See attached image. Files shown are (L-R):
  • luci.20120305.0074
  • luci.20120305.0075
  • luci.20120305.0077
  • Brightest star: BD+32 657 R~9.7 is an M2 star, H=5.036

The data are in H band. The brightest star is
  • Ghosts from a bright star:
    ghosts.jpg

PM: Spectrophotometry tests of the new detector and readout modes Duration: 3h
Responsible: DT, JH
In principle not necessary since RON is known, but since spectroscopy is the dominant observing mode it might be wise to check.
Pre-requisite: spectrophot. std

> Specphot tests of new detector
  • Configure for specphot (wide slit)
  • Preset to any spectrophotometric standard
  • Take data over all major modes (gratings and bandpasses)
  • zJspec at high priority as it is new

Specphot test table
Camera Filter Grating Mask Notes
N1.80 clear K 210_zJHK 3_slit #3 10x30" wide slits

Check for a complete science program (SSC in M51, longslit) Done

20120305: luci.20120305. *

*SSC*
0151 Acquisition-frame
0152 Image slit
0153 Throughslit, OK centered
0154 Offset +30", throughslit OK
0155 Offset, -30", throughslit OK
0156-0165, 1st set of spectra
0166-0175, 2nd set of spectra
0176-0183 Flats
0184-0191 Arcs
*Telluric 1*
0192 Acq. Star 4, in center
0193-0198, Std star obs.
0199-0206, Flats
0207-0214, Arcs
*Telluric 2*
0215 Acq star 1, star off
0216 Check centering, off
0217 check centering, OK
0218-0223 std star obs. fainter than star 4
0224 throughslit, OK
0225-0232 Flats
0233-0240 Arcs

Darks for M51-program 20120306: luci.20120306. *

0006 - 0010 5 images 5 * 2sec integrated, O2DCR
0016-0020 5 images 300sec, normal, MER
0021-0025 5 images 4sec, normal O2DCR
0026-0030 5 images 10sec, normal, O2DCR

LUCI-to-AGw Focus Measurement Accepted as is Duration: 2h
Responsible: DT, JH (+WS)
Pre-requisite: STABLE temps, good (<0.7") seeing

Note: Accepted as is. We have had several hours on sky with ~3 pixel or less FWHM on the N3.75 camera and overall things look good. So there is no evidence of any offset at this time.

Walter, can you outline how you did this?

Note: Can monitor the seeing from the guide cam during measurements to check for any systematic variations. The guide cam will always be in focus once collimation has converged.

> LUCI-to_AGw Focus Measurement

Still to add to the plan?
  • Update/check the collimation model (will happen in course of other tests)
  • Update/check the pointing model (will happen in course of other tests)

-- DavidThompson - 27 Feb 2012

I Attachment Action Size Date Who Comment
20120309.loglog 20120309.log manage 2 K 09 Mar 2012 - 12:32 DavidThompson H band flats/photometry in poor conditions
JHKzeropoints.jpgjpg JHKzeropoints.jpg manage 71 K 06 Mar 2012 - 13:27 DavidThompson JHK Zero Points, March 2012
Repro_dat.tar.gzgz Repro_dat.tar.gz manage 1 K 29 Feb 2012 - 23:53 DavidThompson  
ghosts.jpgjpg ghosts.jpg manage 131 K 06 Mar 2012 - 13:38 DavidThompson Ghosts from a bright star
Topic revision: r32 - 04 Feb 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback